If you can answer any 12 questions from total of 15 questions, how many different ways can you select the questions?

Answer by jim_thompson5910(35256)

C(n,r) = (n!)/(r!(n-r)!)

C(15,12) = (15!)/(12!*(15-12)!)

C(15,12) = (15!)/(12!*3!)

C(15,12) = (15*14*13*12!)/(12!*3!)

C(15,12) = (15*14*13)/(3!)

C(15,12) = (15*14*13)/(3*2*1)

C(15,12) = (2730)/(6)

C(15,12) = 455

So there are 455 ways.